Sonar is a system similar to radar except that sound is used in place of radio waves. Sonar systems

are generally operated under water, where the speed of sound is approximately 1500 m/s. Suppose you are on a ship monitoring the distance to a submarine using sonar, and you find that

the sonar reflections return in 24 s. How far away is the submarine?


1
Expert's answer
2021-01-25T14:03:33-0500

We can find the distance to the submarine, from the formula:


"d=vt=1500\ \dfrac{m}{s}\cdot\dfrac{24\ s}{2}=18000\ m."